OpenStudy (loser66):

Let X and Y have the joint pdf \(f(x,y) = 8x(1-y)\) 0

OpenStudy (loser66):

I stuck at compute \(P(Y<X, X\leq 1/4)\)

OpenStudy (freckles):

\[P(X<Y|X \le \frac{1}{4})=\frac{P(X \le Y,X \le \frac{1}{4})}{P(X \le \frac{1}{4})}\] ok so you said you are having a problem with the numerator \[P(X \le Y,X \le \frac{1}{4})=\int\limits_{0}^{1}\int\limits_{0}^{1-y}8x(1-y) dx d y\] if I remember right

OpenStudy (loser66):

The limits is not good to me.

OpenStudy (freckles):

yes I forgot to include the inequalities

OpenStudy (freckles):

0<y<1 and 0<x<1-y ---- x<=1/4 x<=y ----  ok I understand how they got their setup